OpenStudy (anonymous):

Model the function rule y=-2x-3 with a table of values.

jimthompson5910 (jim_thompson5910):

Pick an x, any x, and plug it into the equation to find y. So say you pick x = 0 y=-2x-3 y=-2(0)-3 ... replace x with 0 y = 0 - 3 y = -3 So when x = 0, y is y = -3 ------------------------------------------------------- Now say we pick x = 1 y=-2x-3 y=-2(1)-3 ... replace x with 1 y = -2 - 3 y = -5 So when x = 1, y is y = -5 ------------------------------------------------------- Now say we pick x = 2 y=-2x-3 y=-2(2)-3 ... replace x with 2 y = -4 - 3 y = -7 So when x = 2, y is y = -7 Which gives us the table  Keep going to fill out the table some more